# CGRclust: Chaos Game Representation for Twin Contrastive Clustering of Unlabelled DNA Sequences

### v1

@@ 1 Background
In 1990, Jeffery introduced a visual representation of DNA sequences, called Chaos Game Representation (CGR) [5]. CGR builds on Barnsley’s 1988 algorithm [6] to map one-dimensional DNA sequences into two-dimensional images, using chaotic dynamics. Several studies [7, 8, 9] have demonstrated that CGRs can act as genomic signatures, defined by Karlin and Burge [10] as numerical quantities that can distinguish closely from distantly related organisms based on DNA sequence identity. The distance between CGRs of DNA sequences can be computed using various metrics, e.g., Euclidean distance, and can then be used for alignment-free comparisons and phylogeny construction to demonstrate evolutionary relationships within a group of organisms. Due to these properties, CGR has been considered a milestone in graphical bioinformatics [11, 12].
@@
Frequency CGR (FCGR) is a quantified variant of CGR: An FCGR at resolution $k$ is a $2^{k}\times 2^{k}$ grayscale image, wherein pixel intensities indicate $k$-mer frequencies. Visual FCGR representations of DNA sequences have been used in many alignment-free genome comparison applications, overcoming the quadratic runtime and scalability problems associated with alignment-based methods [7, 8, 13, 14]. Furthermore, the use of FCGR permits alignment-free genomic sequence comparisons, when used in conjunction with digital signal processing techniques [15, 16] and machine learning methods [17, 18, 19, 20, 21, 22].
@@
FCGR’s ability to convert variable-length sequences into fixed-size dimensions is a key capability for machine learning, especially in DNA classification using convolutional neural networks (CNNs) [23]. In [17], CNNs outperformed Support Vector Machines (SVMs) in classifying FCGR images of bacterial 16S gene sequences for both full-length sequences and 500 bp fragments. In [24], a simple CNN achieved an accuracy of 87% in classifying FCGRs of 660 DNA sequences across eleven genomic datasets. In 2023, Avila et al. effectively classified SARS-CoV-2 DNA sequences into eleven clades using FCGR and CNNs [25], achieving 96.29% accuracy utilizing a ResNet50 neural network [26] and outperforming Covidex [27], a random forest-based clade assignment tool. A hybrid CGR-based approach for detecting COVID-19 was introduced in [28], analyzing both whole and partial genome sequences of 7,951 human coronaviruses using AlexNet, Lasso algorithm, and KNN classifier. In spite of the effectiveness of these DNA classification methods, their reliance on labelled data is a significant limitation which highlights the urgent need for unsupervised algorithms that can perform well without the need of DNA sequence labels.
@@
To address this gap, dense neural networks have been used in conjunction with FCGR for the unsupervised clustering of DNA sequences in large, diverse datasets (up to 9,027 genomes) across different taxonomic levels and genetic distances [29, 30]. However, in both DeLUCS [29] and iDeLUCS [30], the neural networks first flattened each two-dimensional FCGR into a one-dimensional $k$-mer frequency vector. As a result, the features of two-dimensional FCGR images were not fully exploited in these methods. Another approach to clustering unlabelled DNA sequences, MeShClust v3.0 [31], used the mean-shift algorithm for generating pairwise identity scores without alignment. MeShClust v3.0 is built on its predecessors, MeShClust v1.0 [32] (a DNA clustering method) and Identity [33] (a sequence alignment identity score predictor), and can efficiently cluster both long sequences, up to 3.7 million basepairs, and large datasets containing up to a million sequences. MeShClust v3.0 was tested on twenty-seven datasets, including twenty-two synthetic datasets and five real biological datasets, such as the human microbiome and maize transposons. In spite of this progress, in [34] it was shown that DeLUCS, iDeLUCS, and MeShClust v3.0 underperform in clustering astrovirus sequences when compared to $K$-means++, even though they were previously validated on other viral datasets. These limitations highlight the need for the development of more robust approaches that can effectively manage the complexities of genetic diversity of a wide range of genomic datasets.
@@
  • Highly accurate clustering of several current datasets of unlabelled viral whole genomes (Astroviridae family into genera; dengue, HCV, HIV-1 species into virus subtypes), with accuracies ranging from 81.77% to 100%, surpassing the other state-of-the-art clustering methods.
@@ 2.1 Datasets
To comprehensively evaluate the performance of CGRclust in clustering DNA sequences, we strategically selected four groups of datasets, comprising diverse genomic data both real and synthetic. The selection rationale was driven by the need to assess the clustering method across different levels of taxonomy with different degrees of relatedness, genomic conservation, and evolutionary dynamics. The Group 1 dataset includes mitochondrial DNA of fish, while the Group 2 dataset includes viral whole genomes. Additionally, to facilitate direct comparisons with established methodologies, we incorporated datasets previously analyzed in [30] and [31] (Group 3 and Group 4 datasets, respectively).
The Group 1 dataset comprised complete mitochondrial DNA (mtDNA) sequences of Cypriniformes (an order of ray-finned fish). This dataset was retrieved from the National Center for Biotechnology Information (NCBI) on Jan 30, 2024, with a filter selecting mtDNA sequences of length between 4 kbp and 25 kbp. Following the removal of “partial” and “unverified” genomes, 2,688 complete mitochondrial genomes of Cypriniformes were collected. At each taxonomic level, the cluster with the highest number of sequences was selected for the lower taxonomic level clustering task. Due to significant variability and imbalance in the number of available sequences across the four taxonomic levels, sequences from clusters with fewer than 50 sequences were discarded. To address the imbalance, in the first three computational tests (Tests 1-3), we established a threshold based on the minimum number of sequences available in a cluster and randomly selected an equivalent number of sequences from the other clusters. Balancing the clusters was not needed in Test 4, as the dataset was already evenly distributed. Table 1 summarizes the dataset details for the Group 1 dataset (Cypriniformes mtDNA). The selection of this group of datasets was motivated by the conservative nature of mtDNA, which is predominantly coding and thus provides a stable framework for assessing clustering methodologies at multiple taxonomic levels. The uniformity of high conservation over the mtDNA genome compared to the regional variation in sequence conservation of the nuclear genome, coupled with its wide use in phylogenetic studies [37, 38], makes mtDNA data an ideal candidate for initial clustering evaluations.
Table 1: Details of the Group 1 dataset in Tests 1 through 4 (Cypriniformes full mitochondrial genomes). Underlined font indicates the cluster with the highest number of sequences, which was subsequently selected for clustering at a lower taxonomic level. The subfamily Cyprininae includes, e.g., the goldfish (Carassius auratus).
@@
To further demonstrate the effectiveness of CGRclust, we assessed its performance across five viral whole genome datasets in the Group 2 dataset: an updated version of the virus family Astroviridae genomes analyzed in [34] (Test 5) and its balanced version (Test 6), an updated version of whole genomes of dengue virus (Test 7), hepatitis C virus (HCV) (Test 8), and human immunodeficiency virus 1 (HIV-1) (Test 9) previously classified in [39] with supervised machine learning methods. Table 2 outlines the details of the Group 2 dataset. In Test 5, 1,089 complete astrovirus genomes were collected, for taxonomic clustering of the sequences from family to genus level. Test 6 uses a cluster-balanced variant of the astrovirus dataset to address the initial label imbalance, thereby ensuring that the clustering results are not skewed by this disparity. All astrovirus sequences were downloaded from NCBI on April 4, 2024, with a filter selecting genome lengths ranging between 5 kbp and 10 kbp. Furthermore, we addressed the clustering of viral sequences at a lower level of species to subtypes in Tests 7-9. This categorizing which is called viral subtyping is crucial for understanding intraspecific variation, tracking epidemiological trends, and developing targeted treatments or vaccines. The dengue virus sequences used in Test 7 were obtained from https://www.ncbi.nlm.nih.gov/genomes/VirusVariation/Database/nph-select.cgi?taxid=12637 using the query parameters “Nucleotide”, “Full-length sequences only”, and “Collapse identical sequences”, resulting in a dataset of 5,868 sequences. Following cluster balancing, we obtained a dengue dataset comprising 1,628 dengue virus whole genomes spanning four distinct subtypes. The HCV genomes utilized in Test 8 were sourced from the LANL sequence database, accessible at https://hcv.lanl.gov/components/sequence/HCV/search/searchi.html, with the query settings “Excluding recombinants”, “Excluding ‘no genotype‘”, “Genomic region: complete genome”, and “Excluding problematic”, resulting in 3,612 whole HCV genomes. After removing clusters with less than 100 sequences and balancing the dataset, we obtained 950 full HCV genomes spanning five different subtypes. Finally, the HIV-1 genomes in Test 9 were retrieved from the Los Alamos (LANL) sequence database, accessible at https://www.hiv.lanl.gov/components/sequence/HIV/search/search.html with query parameters “virus: HIV-1, genomic region: complete genome, excluding problematic,” which resulted in a dataset comprising 20,525 HIV-1 full genomes. We then removed HIV-1 subtypes with fewer than 100 sequences and balanced the remaining subtypes, thus obtaining a dataset comprising 13,000 HIV-1 whole genome sequences spanning 13 subtypes. The three viral datasets used in Tests 7-9 were downloaded on April 1, 2024. Viral genomes are characterized by higher mutation rates and greater evolutionary diversity compared to the mtDNA, presenting distinct challenges for clustering algorithms. This variability tests the robustness and adaptability of CGRclust under conditions of rapid genomic changes and diverse evolutionary pressures.
Table 2: Details of the Group 2 dataset in Tests 5 through 9 (viral whole genomes).
Test
Taxonomic clustering
(No. of seq. per cluster)
No. of
seq.
Min. seq.
len. (bp)
Avg. seq.
len. (bp)
Max. seq.
len. (bp)
@@
Next, we evaluated the performance of CGRclust on three core datasets used in [30] (Group 3 dataset: mtDNA of Insects, Protists, and Fungi), as well as 12 synthetic DNA datasets analyzed in [31] (Group 4 dataset: synthetic sequences). Including these datasets allowed for direct comparisons with existing studies, providing benchmarks against established clustering methods. The Group 3 dataset is described in Table 3. Note that, given the observed mixed taxonomic levels used in [30] for clustering the Fungi dataset, and the fact that both subphyla “Pezizomycotina” and “Saccharomycotina” belong to phylum Ascomycota, we divided this clustering task into two parts, Tests 12 and 13. The first task (Test 12) involved clustering kingdom Fungi into phyla “Ascomycota” and “Basidiomycota”, while the second task (Test 13) focused on clustering phylum Ascomycota into subphyla “Pezizomycotina” and “Saccharomycotina”. Details about the Group 4 dataset [31] are presented in Table 4. The sequence lengths of six datasets, each beginning with the prefix “Medium-” range between 653 and 2,062 bp, while the other six datasets, prefixed with “Long-”, span from 1,393 to 4,049 bp. The numerical values ranging from 60 to 97 in the dataset labels represent the identity score, a measure of designed relatedness determined by the ratio of identical nucleotides in two sequences relative to the alignment length (including gaps). These synthetic sequences, designed with different sequence lengths and identity score thresholds, evaluate the performance of CGRclust under controlled, and different conditions. For further details on Group 3 and Group 4 datasets, the reader is referred to, [30], and [31], respectively.
Table 3: Details of the Group 3 dataset in Tests 10 through 13 (mtDNA of Insects, Protists, and Fungi, from [33]).
Test
Taxonomic clustering
(No. of seq. per cluster)
No. of
seq.
Min. seq.
len. (bp)
Avg. seq.
len. (bp)
Max. seq.
len. (bp)
@@ 2.2 Method overview
The contrastive clustering method proposed in this paper, CGRclust, utilizes a quantified variant of CGR, a graphical encoding of DNA sequences introduced by Jeffrey in [5]. This quantified DNA encoding, referred to as $FCGR$, represents a DNA sequence at resolution $k$ as a two-dimensional unit square image. In an FCGR, the intensity of each pixel signifies the frequency of a particular $k$-mer in the input DNA sequence [13]. To capture the positional information (location of points) within FCGR images, a CNN model was integrated into the pipeline. CGRclust enhances the clustering performance by leveraging unsupervised contrastive learning. Contrastive learning is a powerful technique that can learn informative representations by comparing how similar or different pairs of examples are, rather than relying solely on raw data or labelled examples [40]. This approach helps the model understand the underlying structures of the data by pulling similar instances (elements of a so-called “positive pair") closer, while pushing dissimilar ones (elements of a so-called “negative pair") farther apart in the representation space. Here, a positive pair is defined as consisting of two “augmented" versions of an input DNA sequence, called mimic sequences. Mimic sequences are generated by the algorithm from an original DNA sequence so as to be similar to the original, or related to it in a meaningful way. In this context, a negative pair is defined as any other pair of sequences in the dataset. The clustering process in this study takes advantage of the concept of twin contrastive learning (TCL) [35, 36], a method that simultaneously optimizes two contrastive learning objectives, one at the instance-level and another at the cluster-level, as detailed below.
@@
To evaluate the quality of the clusters, an additional step, independent from the previous components, is conducted. This step employs the Hungarian algorithm [41] to determine the optimal correspondence between the cluster assignments learned by the CGRclust and the actual taxonomic cluster labels. Subsequently, it evaluates the accuracy of the CGRClust predictions.
2.3 Chaos Game Representation (CGR) of DNA sequences
In the remainder of this paper, the DNA alphabet $\Delta$ is the set $\{A,C,G,T\}$ corresponding to the four different nucleotides, Adenine, Cytosine, Guanine, and Thymine, the building blocks that form a DNA sequence. A CGR square is a square with vertices (corners) in the set $V=\{(1,1),(1,-1),(-1,-1),$ (1,1)}(-1,1)\}( - 1 , 1 ) }. The corners of the CGR square are labelled as follows: the bottom left corner is labelled by $A$, the top left corner is labelled by $C$, the top right corner is labelled by $G$, and the bottom right corner labelled by $T$. Formally, the labelling function $l:\Delta\rightarrow V$ is defined as $l(A)=(-1,-1)$, $l(C)=(-1,1),$ $l(G)=(1,1)$ and $l(T)=(1,-1).$
Definition 1.
Let $s=a_{1}a_{2}\ldots a_{n}$, where $a_{i}\in\Delta$ for all $1\leq i\leq n$, be a DNA sequence of length $n$. A CGR representation $X_{s}$ of the sequence $s$ is the set of points $X_{s}=\{p_{0},p_{1},..,p_{n}\}\in\mathbb{R}^{2}$, situated inside the CGR square, whose coordinates are defined recursively by $p_{0}=(x_{0},y_{0})=(0,0)$ and $p_{i}=\frac{p_{i-1}+l(a_{i})}{2},$ for all $1\leq i\leq n$.
The points plotted in the CGR correspond to nucleotide occurrences in the sequence, and the last plotted point of a CGR could in theory (at infinite resolution) be used to recover the original DNA sequence [42, 43, 44]. Frequency CGR (FCGR) is a quantified variant of CGR: An FCGR of resolution $k$ is a $2^{k}\times 2^{k}$ grayscale image wherein the intensity of each pixel directly corresponds to the frequency of its corresponding $k$-mer. FCGR is a compressed representation of DNA sequences with the compression degree indicated by the resolution $k$. It is obtained by dividing the CGR image into smaller, equally-sized squares and counting the number of points (or the frequency) within each square. Since FCGR smoothes the data in a grayscale image, rather than plotting each point individually, it is inherently less noisy than the traditional CGR. Similar to CGR, FCGR can identify over- and under-representation of patterns (specific arrangements or sequences of nucleotides) in DNA sequences and, thereby, can be used to determine the degree of identity between the DNA sequences of different species [45]. For a formal definition of FCGR at resolution $k$ the reader is referred to Supplementary Material 1.
Figure 2 illustrates some examples of FCGRs at resolution $k=8$ (selected for visualization purposes) of real genomic DNA sequences, side by side with FCGRs of computer-generated DNA sequences.

### v2

@@ 1 Introduction
In 1990, Jeffery introduced Chaos Game Representation (CGR), a method for mapping one-dimensional DNA sequences into two-dimensional space using chaotic dynamics [5, 6]. A CGR maps each DNA sequence to a unique image. The process begins with a unit square whose corners are labelled A, C, G, and T, in a clockwise order starting from the bottom-left corner. The initial point in any CGR plot is the center of this square. To generate the CGR for a specific DNA sequence, the sequence is read from left to right, one nucleotide at a time. For each nucleotide read, a point is plotted midway between the previous point and the corner labelled with that nucleotide. Several studies [4, 7, 8] have demonstrated that CGRs can act as genomic signatures, defined by Karlin and Burge [9] as numerical quantities that can distinguish closely from distantly related organisms based on DNA sequence identity. The distance between CGRs of DNA sequences can be computed using various metrics, e.g., Euclidean distance, and can then be used for alignment-free comparisons and phylogeny construction to demonstrate evolutionary relationships within a group of organisms. Due to these properties, CGR has been considered a milestone in graphical bioinformatics [10, 11].
@@
Frequency CGR (FCGR), a quantified variant of CGR, divides the CGR into smaller squares to calculate and display the frequency of nucleotides within each segment. An FCGR at resolution $k$ creates a $2^{k}\times 2^{k}$ numerical matrix which can be presented as a grayscale image wherein pixel intensities represent $k$-mer frequencies. Consequently, FCGR provides a compressed representation of DNA sequences and facilitates the analysis of distinct genomic signatures across different species. Figure 1 illustrates some examples of FCGRs at resolution $k=8$ (selected for visualization purposes) of real genomic DNA sequences, side by side with FCGRs of computer-generated DNA sequences. FCGR representations of DNA sequences have been used in many alignment-free genome comparison applications, overcoming the quadratic runtime and scalability problems associated with alignment-based methods [12, 7, 13, 14]. The use of FCGR permits alignment-free genomic sequence comparisons, when used in conjunction with digital signal processing techniques [15, 16] and machine learning methods [17, 18, 19, 20, 21, 22].
@@
FCGR’s ability to convert variable-length sequences into fixed-size dimensions is a key capability for machine learning, especially in DNA classification using convolutional neural networks (CNNs) [23]. In a study by Rizzo et al. [17], CNNs outperformed Support Vector Machines (SVMs) in classifying FCGR images of bacterial 16S gene sequences for both full-length sequences and 500 bp fragments. Moreover, Safoury et al. [24] achieved an accuracy of 87% with a simple CNN in classifying FCGRs of 660 DNA sequences across eleven genomic datasets. In 2023, Avila et al. effectively classified SARS-CoV-2 DNA sequences into eleven clades using FCGR and CNNs [25], achieving 96.29% accuracy utilizing a ResNet50 neural network [26] and outperforming Covidex [27], a random forest-based clade assignment tool. Hammad et al. [28] introduced a hybrid CGR-based approach for detecting COVID-19, analyzing both whole and partial genome sequences of 7,951 human coronaviruses using AlexNet, Lasso algorithm, and KNN classifier. In spite of the effectiveness of these DNA classification methods, their reliance on labelled data is a significant limitation which highlights the urgent need for unsupervised algorithms that can perform well without the need of DNA sequence labels.
@@
To address this gap, dense neural networks have been used in conjunction with FCGR for the unsupervised clustering of DNA sequences in large, diverse datasets (up to 9,027 genomes) across different taxonomic levels and genetic distances [29, 30]. However, in both DeLUCS [29] and iDeLUCS [30], the neural networks first flattened each two-dimensional FCGR into a one-dimensional $k$-mer frequency vector. As a result, the features of two-dimensional FCGR images were not fully exploited in these methods. Another approach to clustering unlabelled DNA sequences, MeShClust v3.0 [31], used the mean-shift algorithm for generating pairwise identity scores without alignment. MeShClust v3.0 is built on its predecessors, MeShClust v1.0 [32] (a DNA clustering method) and Identity [33] (a sequence alignment identity score predictor), and can efficiently cluster both long sequences, up to 3.7 million basepairs, and large datasets containing up to a million sequences. MeShClust v3.0 was tested on twenty-seven datasets, including twenty-two synthetic datasets and five real biological datasets, such as the human microbiome and maize transposons. In spite of this progress, DeLUCS, iDeLUCS, and MeShClust v3.0 underperform in clustering astrovirus sequences when compared to $K$-means++ [34], even though they were previously validated on other viral datasets. These limitations highlight the need for the development of more robust approaches that can effectively manage the complexities of genetic diversity of a wide range of genomic datasets.
@@
  • Highly accurate clustering of several current datasets of unlabelled viral whole genomes (Astroviridae family into genera; dengue, HCV, HIV-1 species into virus subtypes), with accuracies ranging from 81.77% to 100% (no classification error), surpassing the other state-of-the-art clustering methods.
@@ 2.1 Datasets
To comprehensively evaluate the performance of CGRclust in clustering DNA sequences, we strategically selected four groups of datasets, comprising diverse genomic data both real and synthetic. The selection rationale was driven by the need to assess the clustering method across different levels of taxonomy with different degrees of relatedness, genomic conservation, and evolutionary dynamics. The Group 1 dataset includes mitochondrial DNA of fish, while the Group 2 dataset includes viral whole genomes. Additionally, to facilitate direct comparisons with established methodologies, we incorporated datasets previously analyzed by Millán et al[30] and by Girgis [31] (Group 3 and Group 4 datasets, respectively). In the following, the test labels are integrally linked with datasets and are used for ease of reference when discussing results.
The Group 1 dataset comprised complete mitochondrial DNA (mtDNA) sequences of Cypriniformes (an order of ray-finned fish). This dataset was retrieved from the National Center for Biotechnology Information (NCBI) on January 30, 2024, with a filter selecting mtDNA sequences of length between 4 kbp and 25 kbp. Following the removal of “partial” and “unverified” genomes, 2,688 complete mitochondrial genomes of Cypriniformes were collected. At each taxonomic level, the cluster with the highest number of sequences was selected for the lower taxonomic level clustering task. Due to significant variability and imbalance in the number of available sequences across the four taxonomic levels, sequences from clusters with fewer than 50 sequences were discarded. To address the imbalance, in the first three computational tests (Tests 1-3), we established a threshold based on the minimum number of sequences available in a cluster and randomly selected an equivalent number of sequences from the other clusters. Balancing the clusters was not needed in Test 4, as the dataset was already evenly distributed. Table 1 summarizes the dataset details for the Group 1 dataset (Cypriniformes mtDNA). The selection of this group of datasets was motivated by the conservative nature of mtDNA, which is predominantly coding and thus provides a stable framework for assessing clustering methodologies at multiple taxonomic levels. The uniformity of high conservation over the mtDNA genome compared to the regional variation in sequence conservation of the nuclear genome, coupled with its wide use in phylogenetic studies [37, 38], makes mtDNA data an ideal candidate for initial clustering evaluations.
Table 1: Details of the Group 1, 2, and 3 dataset in Tests 1 through 13. Underlined font indicates the cluster with the highest number of sequences, which was subsequently selected for clustering at a lower taxonomic level.
Group 1: Cypriniformes Full Mitochondrial Genomes
@@
Group 2: Viral Whole Genomes
@@
Group 3: mtDNA of Insects, Protists, and Fungi (from [30])
@@
To further demonstrate the effectiveness of CGRclust, we assessed its performance across five viral whole genome datasets in the Group 2 dataset: an updated version of the virus family Astroviridae genomes [34] (Test 5) and its balanced version (Test 6), an updated version of whole genomes of dengue virus (Test 7), hepatitis C virus (HCV) (Test 8), and human immunodeficiency virus 1 (HIV-1) (Test 9) previously classified by Solis-Reyes et al. [39] with supervised machine learning methods. Table 1 outlines the details of the Group 2 dataset. In Test 5, 1,089 complete astrovirus genomes were collected, for taxonomic clustering of the sequences from family to genus level. Test 6 uses a cluster-balanced variant of the astrovirus dataset to address the initial label imbalance, thereby ensuring that the clustering results are not skewed by this disparity. All astrovirus sequences were downloaded from NCBI on April 4, 2024, with a filter selecting genome lengths ranging between 5 kbp and 10 kbp. Furthermore, we addressed the clustering of viral sequences at a lower level of species to subtypes in Tests 7-9. This categorizing which is called viral subtyping is crucial for understanding intraspecific variation, tracking epidemiological trends, and developing targeted treatments or vaccines. The dengue virus sequences used in Test 7 were obtained from https://www.ncbi.nlm.nih.gov/genomes/VirusVariation/Database/nph-select.cgi?taxid=12637 using the query parameters “Nucleotide”, “Full-length sequences only”, and “Collapse identical sequences”, resulting in a dataset of 5,868 sequences. Following cluster balancing, we obtained a dengue dataset comprising 1,628 dengue virus whole genomes spanning four distinct subtypes. The HCV genomes utilized in Test 8 were sourced from the LANL sequence database, accessible at https://hcv.lanl.gov/components/sequence/HCV/search/searchi.html, with the query settings “Excluding recombinants”, “Excluding ‘no genotype‘”, “Genomic region: complete genome”, and “Excluding problematic”, resulting in 3,612 whole HCV genomes. After removing clusters with less than 100 sequences and balancing the dataset, we obtained 950 full HCV genomes spanning five different subtypes. Finally, the HIV-1 genomes in Test 9 were retrieved from the Los Alamos (LANL) sequence database, accessible at https://www.hiv.lanl.gov/components/sequence/HIV/search/search.html with query parameters “virus: HIV-1, genomic region: complete genome, excluding problematic,” which resulted in a dataset comprising 20,525 HIV-1 full genomes. We then removed HIV-1 subtypes with fewer than 100 sequences and balanced the remaining subtypes, thus obtaining a dataset comprising 13,000 HIV-1 whole genome sequences spanning 13 subtypes. The three viral datasets used in Tests 7-9 were downloaded on April 1, 2024. Viral genomes are characterized by higher mutation rates and greater evolutionary diversity compared to the mtDNA, presenting distinct challenges for clustering algorithms. This variability tests the robustness and adaptability of CGRclust under conditions of rapid genomic changes and diverse evolutionary pressures.
Next, we evaluated the performance of CGRclust on three core datasets used by Millán et al. [30] (Group 3 dataset: mtDNA of Insects, Protists, and Fungi), as well as 12 synthetic DNA datasets analyzed by Girgis [31] (Group 4 dataset: synthetic sequences). Including these datasets allowed for direct comparisons with existing studies, providing benchmarks against established clustering methods. The Group 3 dataset is described in Table 1. Note that, given the observed mixed taxonomic levels used by Millán et al. [30] for clustering the Fungi dataset, and the fact that both subphyla “Pezizomycotina” and “Saccharomycotina” belong to phylum Ascomycota, we divided this clustering task into two parts, Tests 12 and 13. The first task (Test 12) involved clustering kingdom Fungi into phyla “Ascomycota” and “Basidiomycota”, while the second task (Test 13) focused on clustering phylum Ascomycota into subphyla “Pezizomycotina” and “Saccharomycotina”. Details about the Group 4 dataset [31] are presented in Table 2. The sequence lengths of six datasets, each beginning with the prefix “Medium-” range between 653 and 2,062 bp, while the other six datasets, prefixed with “Long-”, span from 1,393 to 4,049 bp. The numerical values ranging from 60 to 97 in the dataset labels represent the identity score, a measure of designed relatedness determined by the ratio of identical nucleotides in two sequences relative to the alignment length (including gaps). These synthetic sequences, designed with different sequence lengths and identity score thresholds, evaluate the performance of CGRclust under controlled, and different conditions. For further details on Group 3 and Group 4 datasets, the reader is referred to [30] and [31], respectively.
@@ 2.2 Method overview
The contrastive clustering method proposed in this paper, CGRclust, utilizes a quantified variant of CGR, a graphical encoding of DNA sequences introduced by Jeffrey [5]. This quantified DNA encoding, referred to as $FCGR$, represents a DNA sequence at resolution $k$ as a two-dimensional unit square image. In an FCGR, the intensity of each pixel signifies the frequency of a particular $k$-mer in the input DNA sequence [13]. For a formal definition of CGR and FCGR see Supplementary Material 1. To capture the positional information (location of points) within FCGR images, a CNN model was integrated into the pipeline. CGRclust enhances the clustering performance by leveraging unsupervised contrastive learning. Contrastive learning is a powerful technique that can learn informative representations by comparing how similar or different pairs of examples are, rather than relying solely on raw data or labelled examples [40]. This approach helps the model understand the underlying structures of the data by pulling similar instances (elements of a so-called “positive pair") closer, while pushing dissimilar ones (elements of a so-called “negative pair") farther apart in the representation space. Here, a positive pair is defined as consisting of two “augmented" versions of an input DNA sequence, called mimic sequences. Mimic sequences are generated by the algorithm from an original DNA sequence so as to be similar to the original, or related to it in a meaningful way. In this context, a negative pair is defined as any other pair of sequences in the dataset. The clustering process in this study takes advantage of the concept of twin contrastive learning (TCL)  [35, 36], a method that simultaneously optimizes two contrastive learning objectives, one at the instance-level and another at the cluster-level, as detailed below.
@@
To evaluate the quality of the clusters, an additional step, independent from the previous components, is conducted. This step utilizes the Hungarian algorithm [41], a method that effectively pairs elements from two sets to minimize the overall mismatch, to determine the optimal correspondence between the cluster assignments learned by the CGRclust and the actual taxonomic cluster labels. Subsequently, it evaluates the accuracy of the CGRClust predictions.